OpenStudy (anonymous):

Please help me.. I'm so confused on this :(.. A culture started with 2,000 bacteria. After 2 hours, it grew to 2,400 bacteria. Predict how many bacteria will be present after 10 hours. Round your answer to the nearest whole number. Note: When solving for k, round to four decimal places. P = ae^k^t @hartnn

OpenStudy (anonymous):

The equation looks like this: \[\LARGE P = Ae^{kt}\]

OpenStudy (anonymous):

@kelliegirl33

undeadknight26 (undeadknight26):

DO u have to use that equation?...its obvios it increases by 200 each hour.

OpenStudy (anonymous):

Yes, I have to use that equation =/.

undeadknight26 (undeadknight26):

well that stinks...can't help u sorry... @hartnn

OpenStudy (anonymous):

\(P\) is the amount of bacteria right? \(t\) is time. You'd know what e is by calculator.

OpenStudy (anonymous):

@Dakotafox79

OpenStudy (anonymous):

No I don't know how to figure out e via calculator ._. this is confusing me :(.. Sorry I suck at Algebra v.v.. @AnImEfReaK

OpenStudy (anonymous):

wanna do it the quick way?

OpenStudy (anonymous):

If there is a quick way... Then sure why not lol..

OpenStudy (anonymous):

lets forget \( P = Ae^{kt}\) and use only the numbers given to you

OpenStudy (anonymous):

you start with \(2000\) and two hours later you have \(2400\) meaning every two hours in increases by a factor of \(\frac{2400}{2000}=1.2\) since you started with \(2000\) you can model this as \[\large P=2000(1.2)^{\frac{t}{2}}\]

OpenStudy (anonymous):

Why is it just showing "\.[\large P=2000(1.2)^{\frac{t}{2}}.\]" to me??

OpenStudy (anonymous):

you notice i only used the numbers you were given, no logs or anything like that we can do it that way to if you like, it is the same, just takes longer now you want the population at 10 hours put \[\large P=2000(1.2)^{\frac{10}{2}}=2000(1.2)^5\] and use a calculator to get the answer

OpenStudy (anonymous):

refresh your browser

OpenStudy (anonymous):

you are not seeing the latex code for some reason, if you fresh your browser you should see it let me know

OpenStudy (anonymous):

Still shows the same thing ._.

OpenStudy (anonymous):

P = 2000(1.2)^(t/2) is the formula you can use

OpenStudy (anonymous):

then at t = 10 you get P = 2000(1.2)^(10/2)= 2000 (1.2)^5 then use a calculator to find it

OpenStudy (anonymous):

Never mind, I see it now.. Using Google Chrome instead of IE sorry.

OpenStudy (anonymous):

yeah, i.e. is a total waste delete it

OpenStudy (anonymous):

On a school PC can't lol.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@loser of course it works with \(P(2)\) you get \[P(2)=2000\times (1.2)^1=24000\] that is how we got it to begin with

OpenStudy (anonymous):

@Dakotafox79 you good from there? now you need a calculator, but that is all

OpenStudy (anonymous):

How do I do a fraction as a power? How would I input that onto a calculator?

OpenStudy (anonymous):

you don't need a fraction as a power the \(\frac{10}{2}=5\)

OpenStudy (anonymous):

Ohhhh okay. So I just multiply 2,400 to the 5th power correct?

OpenStudy (anonymous):

here lets use this http://www.wolframalpha.com/input/?i=2000%281.2%29^5

OpenStudy (anonymous):

Oh okay I see.

OpenStudy (anonymous):

oh hold on, you are a bit confused still i think formula is \[\large P=2000\times(1.2)^{\frac{t}{2}}\]

OpenStudy (anonymous):

4976.64 is the answer correct?

OpenStudy (anonymous):

\[P(10)=2000\times (1.2)^{\frac{10}{2}}=2000\times (1.2)^5\] exponent first, multiplication last

OpenStudy (anonymous):

yes, but it says round to the nearest whole number

OpenStudy (anonymous):

It says round to the nearest whole number which would be 4977.

OpenStudy (anonymous):

4977

OpenStudy (anonymous):

@Loser66 when you find \(k\) you do some rounding that is why it is not exact

OpenStudy (anonymous):

So that's all it is @satellite73 4977? :O

OpenStudy (anonymous):

yes we can do it the other way if you like

OpenStudy (anonymous):

Nope, that way works for me xD. Much easier.. God you're a hero! A life saver! xD Good job dude! :D Thanks so much :)

OpenStudy (anonymous):

but notice again that i used only the numbers given to you you were given 2000 , 2400 , 2 and from that alone we got the equation

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yw